A projectile is projected with speed 'u' at an angle of 60°  with horizontal from the foot of an inclined plane.If the projectile hits the inclined plane horizontally , the range on inclined plane will be-(in terms of 'u' and 'g')

Grade:11

1 Answers

askIITianexpert IITDelhi
14 years ago

The projectile hits the inclined plane horizontally means at the instant of impact the projectile has only horizontal velocity.

or,it's vertical velocity is zero.

or,it is on it's highest point in the parabolic path.

or,it has gone only half the range for the given angle.

So, height of point of impac on inclined plane=height of projectile=u2Sin2θ/2g=3u2/8g

distance(length)of point of impact on inclined plane=1/2*range of projectile=u2Sinθ.Cosθ/g=√3u2/4g

So,range on inclined plane= [(length)2+(height)2]1/2 =√21u2/8g